For a hypothetical economy, suppose the consumption equation is C = 30 + 0.65Y, where C...

  1. For a hypothetical economy, suppose the consumption equation is C = 30 + 0.65Y, where C is households’ consumption and Y is disposable income. If the disposable income is $600, carefully show your steps and calculate:
    1. The levels of (i) consumption and (ii) saving.
    2. The marginal propensity to (i) consume and (ii) save.
    3. The average propensity to (i) consume and (ii) save.
    4. Interpret your estimates in (c).
Answer #1

i) Consumphan = 420 (= 30 +0.65.1%) = 30+ 0.651 600) 420 ;) Saving 180 = 180 S= y- c = 600-420 = 180 6) i) Maryisal propensit